OpenStudy (anonymous):

using the formula C=2πr, find the circumference of a circle with a diameter of 28 inches.

OpenStudy (nikato):

Given the diameter is 28, can you find the radius? Remember diameter= 2 x radius

OpenStudy (anonymous):

Yes, that is what I remember :)

OpenStudy (nikato):

Ok. So what would be the radius of this circle?

OpenStudy (anonymous):

Just do not know how to set up. Confused:(

OpenStudy (anonymous):

would I set it up something like this: Diameter = (2)(3.14) which becomes -> 28=(2)(3.14)? Am I on the right track? :)

OpenStudy (nikato):

No. All I asked was for the radius. Haha. So Diameter=2xradius. And you know diameter=28 So 28=2 x radius Divide both sides by 2 So Radius=14

OpenStudy (nikato):

Now plug into your circumference formula C= 2(3.14)14

OpenStudy (nikato):

And I'm not sure if you could use it But another version of the formula is C= pi x diameter This would be easier since you were already given the diameter So C=(3.14)28 But you will still get the same answer

OpenStudy (anonymous):

Oh, Thanks. Would never had thought of that! Haha! not good with circles, circumference, etc. Thank you so much ;)

OpenStudy (nikato):

No problem. This is just about plugging numbers into formulas. Nothing much

OpenStudy (jakyfraze):

Everybody the answer is (88) Oh okay. pi= 3.14 x 2 = 6.28. 6.28 x 14 = 87. 92. ANSWER 88 :D right. Round to the nearest 10 which is 88 :)